OpenStudy (anonymous):

A megabyte is 2^20 bytes. A gigabyte is 2^30 bytes. Using the exponent rule for division, find how many megabytes are in a gigabyte.

OpenStudy (anonymous):

that is 2^30/2^20=2^30-20=2^10 times